OpenStudy (anonymous):

Solving Equations Quadratic in Form: x^4-13x^2+36=0

OpenStudy (anonymous):

\[x^{4}-13^{2}+36=0\]

OpenStudy (anonymous):

@jim_thompson5910 can you help me I am stuck on this problem

jimthompson5910 (jim_thompson5910):

Have you tried letting z = x^2 So z^2 = (x^2)^2 = x^4 So z = x^2 and z^2 = x^4 ----------------------------------- This means that x^4-13x^2+36=0 becomes z^2-13z+36=0

jimthompson5910 (jim_thompson5910):

Now ignore everything else for now

jimthompson5910 (jim_thompson5910):

and solve z^2-13z+36=0

OpenStudy (anonymous):

that is where I am at but I can't seem to find anything that adds up to 13 and and 36 (x+)(x-)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how about -9 and -4

OpenStudy (anonymous):

OMG I should take a break cause that was so simple. Maybe not feeling well and math are not good together

jimthompson5910 (jim_thompson5910):

true, but you'll get better at both

OpenStudy (anonymous):

thank u @jim_thompson5910 I can always rely on you

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

keep in mind once you have 2 solutions for z, you use them to find the solutions for x (you'll get 4 solutions total for x)